a boat travels 2 miles upstream and 2 miles downstream in 5 hours. in still water, the boat travels 2 miles/hour.
1. what is speed of river's current?
2. how fast does it travel upstream?
3. how fast does it travel downstream?
4. how long does it take to travel downstream?
5. how long does it take to travel upstream?
Thank you. Gwen

Answer by mananth(16949)

let current speed = x
boat speed =2 mph
against current 2-x
with current x+2
2/(x+2)+2/(2-x) = 5
2(2-x)+2(x+2)=5(4-x^2)
4-2x+2x+4 = 20-5x^2
5x^2=12
x^2= 12/5
x= sqrt(12/5)
upstream speed = 2- 1.55= 0.45 mph
downstream speed = 2+sqrt(12/5)= 3.55 mph
downstream time = 2/3.55= 0.56 hours
upstream time = 2/0.45=4.44
Total time = 5 hours